Question 2: A consumer has 100 dollars to spend on two goods, price of a sweater is $ 10 and price of shoes is $ 20. Assume the consumers budget line has shoes on the vertical axis and sweaters on the horizontal axis;

a. Determine the slope and the intercept of the budget line

b. What will be the effect of the following on the budget line:

Increase in price of sweaters

Increase in income of consumer

If price of a shoes decreases to $ 10 what is the new intercept of the budget line
Answer #1

A.Slope of budget line=-(P1/P2)

P1=Price of sweater

P2=Price of shoes

Slope=-(10/20)=.5

M=Income

X intercept=(M/P1,O)

X intercept=(100/10,0)=(10,0)

Y intercept (0,M/P2)

Y intercept=(0,100/20)=(0,5)

B.1.Increase in price of sweater will reduce the expenditure of sweater.The budget line will shift inward.

2.Increase in income of consumer shifts the budget line to the right, because more of both goods can be consumed.

3.Y intercept=(0,100/10)=(0,10)